OpenStudy (anonymous):

I'm having trouble with Completing the Square in Algebra II... Can someone help me? I have a few practice problems I need help understanding. I'll be posting the problems shortly.

OpenStudy (anonymous):

You must show your work on each of the following questions.

OpenStudy (anonymous):

Ok so completing the square is pretty simple for the x^2+24x+90=0 do you know how we would start that off?

OpenStudy (anonymous):

isolate the x variables

OpenStudy (anonymous):

oh wait... Wouldn't we factor?

OpenStudy (anonymous):

nope you got it right isolate the values

OpenStudy (anonymous):

then take the middle value divide it by 2 then double it (multiply by 2) and square root it and then the solution will be +/-

OpenStudy (anonymous):

its difficult to explain but i hope you get it

OpenStudy (anonymous):

@amistre64 can you help me?

OpenStudy (anonymous):

in COMPLETING THE SQUARE method you just add the "SQUARE OF THE HALF OF THE COEFFICIENT OF THE LINEAR TERM I.E. TERM WITH SINGLE X"

OpenStudy (anonymous):

Okay... Can you help with the first one?

OpenStudy (anonymous):

which one??

OpenStudy (anonymous):

The first one in the picture I posted.

OpenStudy (anonymous):

I assume here that by General Form you mean the Standard Form i.e. f(x) = a(x-h)² + k Since the quadratic is already in general form of f(x)= ax² + bx + c. y = 3x² - 24x + 10 y = 3(x² - 8x) + 10 Now inside the bracket the coefficient of the linear term i.e. 8x is 8 Also half of 8 is 4 So we add as well as subtract square of 4 as below y = 3(x² - 8x + (4)² - (4)² ) + 10 y = 3(x² - 8x + 16 - 16) + 10 y = 3(x² - 8x + 16) - 48 + 10 But (x² - 8x + 16) = (x - 4)² So, y = 3((x - 4)² ) - 48 + 10 y = 3(x - 4)² - 38